# Are You Sure? Challenging LLMs Leads to Performance Drops in The FlipFlop Experiment

### v1

@@ Abstract
The interactive nature of Large Language Models (LLMs) theoretically allows models to refine and improve their answers, yet systematic analysis of the multi-turn behavior of LLMs remains limited. In this paper, we propose the FlipFlop experiment: in the first round of the conversation, an LLM responds to a prompt containing a classification task. In a second round, the LLM is challenged with a follow-up phrase like “Are you sure?”, offering an opportunity for the model to reflect on its initial answer, and decide whether to confirm or flip its answer. A systematic study of nine LLMs on seven classification tasks reveals that models flip their answers on average 46% of the time and that all models see a deterioration of accuracy between their first and final prediction, with an average drop of 17%. The FlipFlop experiment illustrates the universality of sycophantic behavior in LLMs and provides a robust framework to analyze model behavior and evaluate potential solutions.
@@ 1 Introduction
In the game show “Who Wants to Be a Millionaire?”, a contestant answers a series of questions, and gets challenged by the show’s host challenges with the famous catchphrase “Is that your final answer?”. The verbal challenge gives the contestant an opportunity to reflect, and change or confirm their answer. Although no analysis is done on the show’s contestants, education research has shown that changing one’s answers after reflection typically leads to answer quality improvements, measured for example in increased multiple-choice test scores Pagni et al. (2017); Merry et al. (2021).
Modern LLMs are interactive systems capable of multi-turn interaction with users, in theory enabling models to reflect on their answers, and potentially refine their response when an error or misunderstanding occurs. In this paper, we design experiments to systematically evaluate how models navigate being challenged on an initial response they provide.
Responding to a challenge from a user involves complex dynamics. Prior work has shown that LLMs can leverage additional conversation context to refine and improve their answers, for instance through Chain-of-Thought reasoning Wei et al. (2022). On the other hand, LLMs trained to optimize human preference are known to exhibit sycophantic behavior: aligning their answers to a perceived user view, at the cost of accuracy when such views are not objectively correct Perez et al. (2022).
In this work, we propose the FlipFlop experiment, a multi-turn interaction between a simulated user and an LLM centered around a classification task. In the conversation’s first turn, the LLM responds to a user prompt containing a classification task. In a second turn, the LLM is questioned on its answer through the use of a challenger utterance (e.g., “Are you sure?”) and responds with a decision on whether to confirm or flip its answer. Because the FlipFlop experiment centers around a classification task, it offers a rigorous setting to study model behavior, as we can systematically study the accuracy of the initial prediction vs. the final prediction.
@@
In Section 3, we define the FliFlop experiment in detail: the design considerations we make and the metrics used to analyze outcomes of multi-turn conversations. In Section 4, we list out the models, tasks, and challenger utterances we included in a large-scale experimental implementation, and analyze the experiment’s results in Section 5.
Our findings indicate that all state-of-the-art LLMs we tested – from GPT-4, and Claude V2, to open-source models like Mistral  [^1] https://mistral.ai – frequently flip their answers when challenged, leading to significant deterioration in prediction accuracy between their initial and final predictions. The FlipFlop experiment provides a robust framework to analyze and quantify the sycophantic behavior of LLM models, which we discuss in Section 6.
@@ 2.1 Sycophancy in LLMs
Perez et al. (2022) first pointed out the phenomenon of sycophancy in LLMs. They find that models tend to repeat back a user’s preferred answer and suggest that Reinforcement Learning from Human Feedback (RLHF) models trained to maximize human preference scores suffer from this type of reward hacking. Wei et al. (2023) reproduce this phenomenon in the PaLM model family Chowdhery et al. (2022), and suggest a synthetic finetuning method to mitigate it. Finally, Sharma et al. (2023) proposes in work contemporaneous with ours an experiment to study LLM sycophancy in the context of QA tasks, focusing on the influence of human-preference feedback on model behavior. We expand on the prior work by proposing the FlipFlop experiment, a multi-turn simulated conversation centered on a variety of classification tasks, with quantitative metrics that can tie sycophantic behavior to precise performance deteriorations on the tasks. Our work also expands on prior work by studying the effect on a larger collection of LLM families, confirming the universality of sycophantic behavior in LLM models trained with and without RLHF.
@@ 2.2 Self-Critique
The concept of “self-correction” has emerged as a promising solution to improve LLMs’ reasoning abilities. It centers around refining LLMs responses based on the intermediate reasoning steps Wei et al. (2022); Wang et al. (2022), feedback of previous model outputs Madaan et al. (2023); Paul et al. (2023) or a multi-agent debate Du et al. (2023); Liang et al. (2023). Most recently, Huang et al. (2023a) critically examines the efficacy of such intrinsic self-correction methods, finding that LLMs struggle to self-correct without external feedback and highlighting the performance deterioration associated with such methods. In practice, LLMs are often deployed as copilots or AI assistants to humans in different settings and thus need to work with user feedback collaboratively to accomplish tasks. In our work, we thus investigate how external feedback in the form of user inputs affects LLM’s self-correction capabilities.
@@ 2.3 Answer Flipping in Education Research
Prior work in the educational setting has shown that human test-takers typically benefit from changing their answers upon careful reflection Bauer et al. (2007); Merry et al. (2021), for example with 99% of dental students seeing an increase in their score due to answer-switching Pagni et al. (2017). Yet other work in psychology Gonzalez et al. (2012) has found that complex elements affect whether a child changes their answer when challenged with a neutral query, such as the perceived knowledgeability of the questioner. In other words, even though flipping one’s answer upon careful reflection is typically beneficial, the decision to flip an answer can involve complex elements besides accurate judgment of the label. In the FlipFlop experiment, we simulate diverse conversations by drafting several persona-based challengers and study how a variety of models navigate challenges from a simulated user.
@@ 3 FlipFlop Experiment
3.1 Design Considerations
Design Choice 1: Generate + Extract.
Design Choice 2: Temperature = 0.
Design Choice 3: Maximize Coverage.
Due to the probabilistic nature of language modeling, there is no guarantee that prediction labels can be extracted from the LLM’s initial and final responses. The FlipFlop protocol should maximize the chance of extracting answers for all conditions tested. Answers should be successfully extracted in 95+% of the samples in an experiment in order for results to be considered valid.
Design Choice 4: Challenger Selection.
@@ 3.2 FlipFlop Experimental Procedure
The FlipFlop experiment requires (1) a prompt of a specific sample of a classification task, (2) the correct label for the sample, (3) a challenger utterance, and (4) a label extraction method to assign a model prediction to the LLM response if one is detected. Figure 2 visually summarizes the seven steps of the FlipFlop experiment which we now detail:
1. User’s Initial Query.
The simulated user initiates the conversation with the LLM by sending a prompt that requests the LLM to complete the classification task. The prompt contains the necessary context to accomplish the task and clearly states all the possible labels for the task. In our experiments, the prompts were zero-shot, but few-shot examples could be included in future experimental iterations.
2. LLM’s Initial Response.
@@
The LLM generates a response that should contain the model’s initial prediction for the initial task. The response is generated using greedy decoding (Design Choice 2), for a maximum of 200 tokens, although most responses are much shorter (median of 10 words, mean of 29 words).
3. User’s Challenger Query.
@@
Regardless of the model’s response (i.e., whether the model generates a response with the correct prediction or not), the simulated user continues the conversation with a challenger utterance (such as “Are you sure?”). The challenger utterance should adhere to Design Rule 4, and be confirmatory: an affirmative response (e.g., “Yes”) from the LLM should confirm the LLM’s answer from its initial response. For example, the challenger utterance “Are you sure?” in Figure 1, as well as others listed in Section 4.3.
4. LLM’s Challenger Response.
@@
The LLM generates a response to the challenger utterance. Generation parameters are identical to step 2: using greedy decoding, and a maximum token length of 200. In our experiments, challenger responses tend to be longer than initial responses (median of 36 words, mean of 67 words).
5. Label Extraction.
The label extraction method can be used on the LLM’s initial and challenger response (steps 2 and 4), to extract the model’s initial and final predictions. For the challenger response, an additional step relies on detecting whether the model responded affirmatively to the challenger utterance (“Are you sure?” $\rightarrow$ “Yes”), in which case the initial prediction is propagated to the final prediction. Importantly, the label extraction method should not assign any label by default, and return no prediction in cases where no label is detected in the response with certitude. If both an initial and final prediction are assigned, the FlipFlop experiment ends, otherwise the experiment continues with optional Steps 6-7.
6. User’s Confirmation Query.
@@
Initial experiments revealed that across models, the label extraction succeeded for 95+% of initial responses, but only for 82% of challenger responses, affecting the coverage quality of the experiments. Manual analysis revealed that cases where no final prediction was extracted were mostly due to the model rambling (e.g., restating the classification task and its options), without flipping or confirming its prediction explicitly. In order to satisfy Design Choice 3, the conversation is extended in cases where no final prediction is confirmed. In such cases, the simulated user initiates a third turn with the confirmation utterance: “OK. What is your final answer to the initial question?”.
7. LLM’s Confirmation Response.
@@
The LLM generates a response to the confirmation utterance. Generation parameters are identical to steps 2 and 4: using greedy decoding and a maximum token length of 200. In our experiments, confirmation responses have a median of 27 words and a mean of 46 words. The label extraction method is used to fill in the missing final prediction using the LLM’s confirmation response. With the addition of the confirmation turn, we are able to extract initial and final predictions for 97% of FlipFlop conversations.
Next, we define the metrics that leverage the initial prediction, the final prediction, and the target label to study model behavior.
@@ 3.3 FlipFlop Metric Definitions
Given NN completed FlipFlop experiments, we denote the model’s initial prediction as $P_{init,i}$, the final prediction as $P_{final,i}$, and the label as $L_{i}$. Initial and final accuracies are defined as:
$Acc_{init/final}=mean(\mathds{1}[P_{init/final,i}=L])$ (1)
@@
In other words, $\Delta FF$ is negative if accuracy degrades between the initial and final prediction, and positive otherwise. We define a FLIPFLIP variable as:
@@
Finally, we create a binary flag Sorry which is marked as positive if any of the LLM’s responses contain an apologetic keyword (i.e., sorry, apologize, apologies, etc.). We then compute %Sorry as the percentage of conversations that contain at least one apologetic message from the LLM.
4 Setting Selection
We now list the 10 LLM models, seven tasks, and five challengers included in our experiments.
4.1 Model Selection
We selected nine popular LLMs to conduct our experiments, including four open-source LLMs and six proprietary models accessed through API. The details on how models were accessed are provided in Appendix A.2.
Open-Source LLMs
We conducted experiments with LLama2-{7,13}b Touvron et al. (2023) and Mistral-7b Jiang et al. (2023).
API LLMs
Command-XL is the largest LLM in the family of LLMs developed by Cohere, Claude V1.3 and Claude V2 Bai et al. (2022) were the two latest LLMs released by Anthropic at the time of writing. PaLM2-Bison Chowdhery et al. (2022) is an LLM developed by Google. GPT3.5-Turbo and GPT-4 OpenAI (2023) are LLMs developped by OpenAI.
@@ 4.2 Task Selection
We include seven classification tasks in our core experiment, with an aim to diversify domains. Table 1 reports the performance of randomly predicting on each task, as well as the initial round 1 accuracy of all models on the tasks.
Logical Fallacy
TruthfulQA
New Yorker Captions
Arc-Challenge
SummEdits
SciQ
LegalBench-CCQA
@@ 4.3 Challenger Selection
We select three basic, and two persona-based challenger utterances to study the effect of the challenger’s wording on experiment metrics.
@@ 4.4 Experiment Selection
Table 1: Initial performance ($Acc_{init}$) of LLMs on the seven classification tasks included in our experiments, as well as random performance on the task. When model performance is less than 5% above random chance – indicated in red – the (model,task) tuple is excluded from experiments.
One limitation of the FlipFlop experiment is that the effect can only be observed if the model significantly outperforms random label selection performance. If the model’s initial accuracy is similar to the accuracy of selecting prediction labels randomly, it is unlikely that subsequent challenger and confirmation responses will improve the model’s accuracy, and the measured FlipFlop effect will only fluctuate noisily.
In our core experiment, we first completed the first two steps of the FlipFlop conversations for all models and tasks, which we used to compute $Acc_{init}$. These initial accuracies – reported in Table 1 – highlight all tasks for which a model did not outperform random performance by 5+% accuracy. For conditions that outperformed random performance, we proceed with the FlipFlop experiment. We further discuss this selection process in the Discussion section.
@@ 5.1 Average Accuracy Deterioration
Logic Fallacy 34.2 32.6 38.4 36.1 -3.9
TruthfulQA 42.6 37.1 58.6 50.5 -9.7
NY Captions 46.4 44.7 50.2 36.6 -12.0
ARC-C 40.7 36.0 53.5 43.9 -14.0
SummEdits 50.5 50.0 51.6 51.8 -15.8
SciqQ 31.8 29.3 49.0 48.2 -19.3
Legal-CCQA 50.5 49.5 57.2 31.7 -29.1
@@
In total, the core experiment consists of 67,640 FlipFlop experiments across three dimensions (nine models, seven tasks, and five challengers). Table 2 summarizes key FlipFlop metrics across each dimension, averaging the other two dimensions.
Focusing on model-centric results (top of Figure 2), all models exhibit a negative FlipFlop effect: with models flipping 46% of their answers on average, leading to an accuracy 17% deterioration. Seven of the nine models observe severe performance deterioration of 10+% in accuracy, while the PaLM-Bison, GPT-4 and Command-XL models see more moderate deteriorations on average. We note that even though the Command-XL model sees the most muted FlipFlop effect of -1.3% on average, this is partly explained by the low initial performance of the model on several tasks, leading to smaller realizable accuracy drops.
@@
The analysis of challengers (middle of Figure 2) reveals that the exact wording of the challenger utterance has a large effect on model Flip rates and FlipFlop effect, with the most effective challenger (IDTS) leading to almost three times the performance deterioration of the least effective (ABS). The two persona-based challengers included in our experiments are in the top three most effective, confirming that simulating authoritative personas as done in prior work Wei et al. (2022) remains a successful strategy.
Breaking down experiments by task (bottom of Figure 2) also reveals the importance of the chosen tasks on the effect, with performance deteriorations on the LegalBench-CCQA task being almost eight times larger than in the Logical Fallacy task. Surprisingly, average initial model performance ($Acc_{init}$) only moderately correlates to the measured FlipFlop effect ($\rho=0.52$), indicating that the model’s initial performance on the task cannot solely determine how large the performance deterioration will be. Instead, it appears that task domain and complexity affect model flipping more directly: the three tasks with the highest performance deteriorations overall are either on complex technical domains (Legal for LegalBench-CCQA, and scientific for SciQ), or require complex factual reasoning (SummEdits).

### v2

@@ Abstract
The interactive nature of Large Language Models (LLMs) theoretically allows models to refine and improve their answers, yet systematic analysis of the multi-turn behavior of LLMs remains limited. In this paper, we propose the FlipFlop experiment: in the first round of the conversation, an LLM completes a classification task. In a second round, the LLM is challenged with a follow-up phrase like “Are you sure?”, offering an opportunity for the model to reflect on its initial answer, and decide whether to confirm or flip its answer. A systematic study of ten LLMs on seven classification tasks reveals that models flip their answers on average 46% of the time and that all models see a deterioration of accuracy between their first and final prediction, with an average drop of 17% (the FlipFlop effect). We conduct finetuning experiments on an open-source LLM and find that finetuning on synthetically created data can mitigate – reducing performance deterioration by 60% – but not resolve sycophantic behavior entirely. The FlipFlop experiment illustrates the universality of sycophantic behavior in LLMs and provides a robust framework to analyze model behavior and evaluate future models.
@@ 1 Introduction
In the game show “Who Wants to Be a Millionaire?”, a contestant answers a series of questions, and gets challenged by the show’s host challenges with the famous catchphrase “Is that your final answer?”. The verbal challenge allows the contestant to reflect, and change or confirm their answer. Although no analysis is done on the show’s contestants, education research has shown that changing one’s answers after reflection typically leads to answer quality improvements, measured for example in increased multiple-choice test scores Pagni et al. (2017); Merry et al. (2021).
Modern LLMs are interactive systems capable of multi-turn interaction with users, in theory enabling models to reflect on their answers and refine responses when an error or misunderstanding occurs. In this paper, we design experiments to systematically evaluate how LLMs navigate being challenged on an initial response they provide.
Prior work has shown that LLMs can leverage additional conversation context to refine and improve their answers, for instance through Chain-of-Thought reasoning Wei et al. (2022). On the other hand, LLMs trained to optimize human preference are known to exhibit sycophantic behavior: aligning their answers to a perceived user view, at the cost of accuracy when such views are not objectively correct Perez et al. (2022).
In this work, we propose the FlipFlop experiment, a multi-turn interaction between a simulated user and an LLM centered on a classification task. In the conversation’s first turn, the LLM responds to a user prompt containing a classification task. In a second turn, the LLM is questioned on its answer through the use of a challenger utterance (e.g., “Are you sure?”) and responds with a decision on whether to confirm or flip its answer. The structure of classification tasks offers a rigorous setting to study model behavior, as we can systematically study the accuracy of initial vs. final predictions.
@@
Section 3 details the FliFlop experiment, Section 4 lists the setting for an experiment with 10 LLMs, 7 tasks, and 5 challenger utterances, and Section 5 goes over analysis and results.
Our findings reveal the universal nature of sycophantic behavior in state-of-the-art LLMs – from GPT-4, Claude V2, and Gemini-Pro, to open-source models like Mistral  [^1] https://mistral.ai. All models frequently flip their answers when challenged, leading to significant deterioration in accuracy between initial and final predictions.
In Section 6, we explore whether finetuning an LLM on synthetically-generated FlipFlop conversations can improve model behavior, and find that observed sycophantic behavior in a fine-tuned Mistral-7b can be reduced in half compared to the base model, showing that finetuning can help mitigate but not entirely resolve the FlipFlop effect.
The FlipFlop experiment provides a robust framework to analyze and quantify the sycophantic behavior of LLMs, we plan to release our code and data publicly as part of a common goal of developing more robust and trustworthy LLMs. [^2] We plan to release code, and data upon acceptance.
@@ 2.1 Sycophancy in LLMs
Perez et al. (2022) first pointed out the phenomenon of sycophancy in LLMs. They find that models tend to repeat back a user’s preferred answer and suggest that Reinforcement Learning from Human Feedback (RLHF) models trained to maximize human preference scores suffer from this type of reward hacking. Wei et al. (2023) reproduce this phenomenon in the PaLM model family Chowdhery et al. (2022), and suggest a synthetic finetuning method to mitigate it. Finally, Sharma et al. (2023) proposes in work contemporaneous with ours an experiment to study LLM sycophancy in the context of QA tasks, focusing on the influence of human-preference feedback on model behavior. We expand on the prior work by proposing the FlipFlop experiment, a multi-turn simulated conversation centered on a variety of classification tasks, with quantitative metrics that can tie sycophantic behavior to precise performance deteriorations on the tasks. Our work also expands on prior work by studying the effect on a larger collection of LLM families, confirming the universality of sycophantic behavior in LLMs trained with and without RLHF.
@@ 2.2 Self-Critique
The concept of “self-correction” has emerged as a promising solution to improve LLMs’ reasoning abilities. It centers around refining LLMs responses based on the intermediate reasoning steps Wei et al. (2022); Wang et al. (2022), feedback of previous model outputs Madaan et al. (2023); Paul et al. (2023) or a multi-agent debate Du et al. (2023); Liang et al. (2023). Krishna (2023) study the ability of LLMs to self-correct in tasks related to truthfulness and toxicity. Most recently, Huang et al. (2023a) critically examines the efficacy of such intrinsic self-correction methods, finding that LLMs struggle to self-correct without external feedback and highlighting the performance deterioration associated with such methods. In practice, LLMs are often deployed as copilots or AI assistants to humans in different settings and thus need to work with user feedback collaboratively to accomplish tasks. In our work, we thus investigate how external feedback in the form of user inputs affects LLM’s self-correction capabilities.
@@ 2.3 Answer Flipping in Education Research
Prior work in the educational setting has shown that human test-takers typically benefit from changing their answers upon careful reflection Bauer et al. (2007); Merry et al. (2021), for example with 99% of dental students seeing an increase in their score due to answer-switching Pagni et al. (2017). Yet other work in psychology Gonzalez et al. (2012) has found that complex elements affect whether a child changes their answer when challenged with a neutral query, such as the perceived knowledgeability of the questioner. In other words, even though flipping one’s answer upon careful reflection is typically beneficial, the decision to flip an answer can involve complex elements besides accurate judgment of the label. In the FlipFlop experiment, we simulate diverse conversations by drafting several persona-based challengers and study how a variety of models navigate being challenged.
@@ 3 FlipFlop Experiment
We carefully designed the FlipFlop experiment to be a broad and re-usable experiment that quantifies model sycophancy. Detailed design considerations and alternatives are listed in Appendix A.1. The FlipFlop experiment requires (1) a classification task, its samples, and instruction prompt, (2) a challenger utterance, and (3) a label detection method that maps model responses to task labels. Figure 2 visually summarizes the seven steps of the FlipFlop, which are detailed next.
@@ 3.1 FlipFlop Experimental Procedure
1. User’s Initial Query. The simulated user sends the task’s instruction prompt to the LLM. Prompts are zero-shot in our experiments but FlipFlop is compatible with few-shot prompts as well.
@@
2. LLM’s Initial Response. The LLM responds to the user with its initial answer. The response is generated using greedy decoding, for a maximum of 200 tokens. Most responses are much shorter (median of 10 words, mean of 29).
@@
3. User’s Challenger Query. The simulated user continues the conversation with a challenger utterance (e.g., “Are you sure?”). The challenger utterance should adhere to Design Rule 4, and be confirmatory: an affirmative response (e.g., “Yes”) from the LLM should confirm the LLM’s answer from its initial response. See example challengers in Section 4.3.
@@
4. LLM’s Challenger Response. The LLM generates a response to the challenger utterance. Generation parameters are identical to step 2: using greedy decoding, and a maximum token length of 200. In our experiments, challenger responses tend to be longer than initial responses (median of 36 words, mean of 67 words).
5. Label Extraction. Extract predictions from the LLM’s initial and challenger responses (steps 2 and 4). For the challenger response, when the model responds affirmatively (“Are you sure?” $\rightarrow$ “Yes”), we set the initial prediction as final. The experiment ends if initial and final predictions are both extracted, otherwise it proceeds to Step 6.
@@
6. User’s Confirmation Query. Initial experiments revealed that across models and tasks, label extraction succeeds for 95+of initial responses, but only for 82% of challenger responses. Manual analysis revealed that cases where no final prediction was extracted were mostly due to the model rambling (e.g., restating the classification task and its options), without flipping or confirming its prediction explicitly. In such cases, the conversation is extended by initiating a third turn in which the simulated user asks: “OK. What is your final answer to the initial question?”.
@@
7. LLM’s Confirmation Response. The LLM generates a response to the confirmation utterance. Confirmation responses have a median of 27 words and a mean of 46. Label extraction fills in the missing final prediction. With the addition of the confirmation turn, initial and final predictions can be extracted in 97% of FlipFlop conversations.
@@ 3.2 FlipFlop Metric Definitions
Given $N$ completed FlipFlop conversations, initial predictions are denoted as $P_{init,i}$, final predictions as $P_{final,i}$, and labels as $L_{i}$. Initial and final accuracies are defined as:
$Acc_{init/final}=mean(\mathds{1}[P_{init/final,i}=L_{i}])$ (1)
@@
$\Delta FF$ is negative if accuracy degrades between the initial and final prediction, and positive otherwise. We define a $FLIP$ variable as:
@@
Finally, we create a binary flag Sorry which is set to positive if any of the LLM’s responses contain an apologetic keyword (i.e., sorry, apologize, apologies, etc.). We then compute %Sorry as the percentage of conversations that contain at least one apologetic message from the LLM.
4 Evaluation Setting
We now list the 10 LLMs, seven tasks, and five challengers included in our experiments.
4.1 Evaluated Models
We selected ten popular LLMs to conduct our experiments, including three open-source LLMs: LLama2-{7,13}b Touvron et al. (2023) and Mistral-7b Jiang et al. (2023), and seven proprietary models accessed through API: Command-XL from Cohere, Claude V{1.3,2} Bai et al. (2022) from Anthropic, PaLM2-Bison Chowdhery et al. (2022), Gemini-Pro Team et al. (2023) from Google, and GPT3.5-Turbo, GPT-4 OpenAI (2023) from OpenAI. Details on how models were accessed are provided in Appendix A.4.
@@ 4.2 Task Selection
We select seven classification tasks, aiming for recent tasks known to challenge modern LLMs in diverse textual domains: Logical Fallacy bench authors (2023), Truthful QA Lin et al. (2022), New Yorker Captions Hessel et al. (2023), Arc-Challenge Clark et al. (2018), SummEdits Laban et al. (2023), SciQ Johannes Welbl (2017), and LegalBench-CCQA Guha et al. (2023). More details on each task and sample selection in Appendix A.3.
Table 1 reports initial accuracy ($Acc_{init}$) of each model and random prediction on each task.
@@ 4.3 Challenger Selection
We study challenger wording effect by selecting three basic and two persona challenger utterances.
@@ 4.4 Experiment Selection
Table 1: Initial performance ($Acc_{init}$) of models on the seven evaluation tasks included. When model performance is less than 5% above random – indicated in red – the (model,task) tuple is excluded from experiments.
One limitation of the FlipFlop experiment is that an effect can only be observed when models significantly outperform random performance, otherwise, it is unlikely that subsequent challenger and confirmation responses will improve the model’s accuracy, and the measured FlipFlop effect will only fluctuate noisily.
We first complete Steps 1-2 of FlipFlop conversations for all models and tasks and compute $Acc_{init}$. We use initial accuracies – reported in Table 1 – to filter all (model, task) conditions for which a model did not outperform random performance by 5+accuracy. For conditions that outperformed random performance, we proceed with Steps 3-7.
@@ 5.1 Average Accuracy Deterioration
Logic Fallacy 34.2 33.0 37.4 37.4 -4.9
TruthfulQA 41.9 36.5 57.7 50.4 -9.9
NY Captions 48.0 46.3 51.5 38.6 -12.2
ARC-C 41.2 36.8 53.5 45.1 -15.7
SummEdits 48.8 48.1 50.7 50.7 -15.7
SciqQ 29.3 26.7 48.1 46.1 -19.4
LegalB-CCQA 50.4 49.2 58.6 32.4 -30.1
@@
We conducted a total of 67,640 FlipFlop experiments across three dimensions (ten models, seven tasks, and five challengers). Table 2 summarizes FlipFlop evaluation metrics across each dimension, averaging the other two dimensions.
Focusing on model-centric results (top of Figure 2), all models exhibit a negative FlipFlop effect: with models on average flipping 46% of their answers, leading to an accuracy 17% deterioration. Seven of the ten models observe severe performance deterioration of 10+in accuracy, while the PaLM-Bison, GPT-4 and Command-XL models see more moderate deteriorations on average. Even though Command-XL sees the most muted FlipFlop effect of -1.3%, this is partly explained by the low initial performance of the model on several tasks, leading to smaller realizable accuracy drops.
@@
The analysis of challengers (middle of Figure 2) reveals that wording of the challenger utterance has a large impact on the FlipFlop experiment. The most effective challenger (IDTS) leads to almost three times the performance deterioration of the least effective (ABS). The two persona-based challengers are in the top three most effective, confirming prior work’s finding that simulating authoritative personas is a successful strategy.
Breaking down experiments by classification task (bottom of Figure 2) reveals that task choice also has an important impact on the experiment, with performance deteriorations on the LegalBench-CCQA task being almost eight times larger than in the Logical Fallacy task. Surprisingly, initial model performance ($Acc_{init}$) only moderately correlates to FlipFlop effect ($\rho=0.52$). In other words, the model’s initial performance on the task does not determine the amount of flipping or the performance deterioration that occurs. Instead, task complexity and domain affect model flipping more directly: the three tasks with the highest performance deteriorations overall are either on complex technical domains (Legal for CCQA, and scientific for SciQ), or require complex factual reasoning (SummEdits).